# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

### v1

@@ 3.4 Dimension Reduction of the Future Trajectory Parameterization
To summarize the equations, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ can be related to $\eta_{1}(k)$ and $\eta_{2}(k)$ using (17) and (39) (via $x_{np}(k)$). Substituting these into (9): The past part is $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{Z}_{p}\mathbf{g}_{f}(k)=0$ and (16), then $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$. The future part is $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ from (17) and $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=L_{f_{1}}\eta_{2}(k)$ from (39). So, $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k)$. By defining $S=\mathbf{Z}_{f}V_{1}$, the past and future input/output data can be written in terms of intermediate variables $\eta_{1}(k)$ and $\eta_{2}(k)$ as follows:
$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix},$ (40)
where $L_{1}$ is partitioned as $\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}$ (implicitly, so $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$ from the top block-row of the matrix equation if zero matrix has appropriate dimensions). And we partition $S$ and $L_{f_{1}}$ matrices via:
$S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$ (41)
@@ 3.5 Optimal Multi-Step Predictor
We now apply the parameterization in (40) to a $N$-step ahead prediction of the future output. Optimality in this context refers to finding the minimum variance unbiased predictor of the future output. Using the lower block of (40), we have the following relation for future trajectories:
$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}S_{u}&L_{f_{u}}\\ S_{y}&L_{f_{y}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$ (42)
where it can be expanded for $\mathbf{u}_{N}(k)$ and $\mathbf{y}_{N}(k)$ as:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (43)
To derive the future control input $\mathbf{u}_{N}(k)$, we substitute the BLUE estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (43), yielding:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (44)
Assuming $L_{f_{u}}$ is invertible, we solve for $\eta_{2}(k)$ from the first equation in (44):
@@
Substituting (45) into the expression for $\mathbf{y}_{N}(k)$ in (44), we obtain the predictor for $\mathbf{y}_{N}(k)$ which we denote $\hat{\mathbf{y}}_{N}(k)$:
@@
Using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}{\mathbf{z}}_{\text{ini},m}(k)$:
$\begin{split}\hat{\mathbf{y}}_{N}(k)&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (47)
Define:
$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})L_{1_{\Sigma_{\zeta}}}^{\dagger},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (48)
so that:
@@
To ensure the mathematical rigor of the NTDPC problem, we introduce the following assumptions and lemma.
@@
The matrix $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is invertible, ensuring that the future input $\mathbf{u}_{N}(k)$ uniquely determines $\eta_{2}(k)$ for a given $\hat{\eta}_{1}(k)$.
@@
The noise $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian with known positive definite covariance $\mathbf{\Sigma}_{\zeta}$. The SVD separates signal and noise such that $\Sigma_{1}$ in (18) captures the dominant system dynamics.
Assumption 3.4.
The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex and non-empty.
@@
Under Assumptions 3.2 and 3.3, the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$ is an unbiased estimator of the true future output $\mathbf{y}_{N}(k)$ that would result from $\mathbf{u}_{N}(k)$ and the true initial state encoded by $\eta_{1}(k)$, i.e., $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]=S_{y}\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$. The matrices $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ can be computed offline.
@@
From (49), $\hat{\mathbf{y}}_{N}(k)$ is derived using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. By Theorem 3.1, $\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]=\eta_{1}(k)$. The true output corresponding to $\eta_{1}(k)$ and a chosen $\mathbf{u}_{N}(k)$ (which defines $\eta_{2}(k)$) is given by the second line of (43): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$. With $\eta_{2}(k)=L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$, this becomes $\mathbf{y}_{N}^{\text{true}}(k)=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)$. Taking the expectation of the predictor (47):
$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (50)
Thus, the predictor is unbiased for the true output that would arise from $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$. The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ follows from their definitions in (48), as they depend on SVD results and system parameters, not online data (except $\mathbf{z}_{\text{ini},m}(k)$ which is an input to the predictor). ∎
We now formulate the NTDPC problem.
Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC solves the following optimization problem at each time step $k$:
@@
s.t. $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (51)
$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\hat{\mathbf{y}}_{N}(k)\in\mathcal{Y},$
where $\mathbf{u}_{N}(k)\in\mathbb{R}^{n_{u}N}$ is the input sequence over horizon $N$, $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the predicted output sequence in the cost, $\hat{\mathbf{y}}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the constrained output sequence (related to $\mathbf{y}_{N}(k)$ through the model), $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable, $\mathbf{z}_{\text{ini},m}(k)=\text{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini},m}(k))\in\mathbb{R}^{(n_{u}+n_{y})T_{\text{ini}}}$ contains past inputs and noisy outputs ($\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$), $\mathbf{P}_{1},\mathbf{P}_{2}$ are SVD-derived predictor matrices, $Q,R,\Lambda_{y}\succ 0$ are weighting matrices, and $\mathcal{U},\mathcal{Y}$ are convex constraint sets. Algorithm 1 summarizes the main steps of the NTDPC approach.
The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ reduces online complexity, and the absence of regularization parameters simplifies tuning compared to DeePC, $\gamma$-DDPC, and GDPC.
@@
When calculating the SVD factorization for matrices $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$ in (12) and (38) in noisy case, it is recommended to normalize the rows of these matrices first, to avoid giving higher priority to input or output. For this purpose, the input and output signals need to be scaled at the first step as $\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k)$ where $M_{u}=\text{diag}(m_{u_{1}},\cdots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\cdots,m_{y_{n_{y}}})$ in which $m_{u_{i}}$ and $m_{y_{i}}$ shows the range of $\mathbf{u}_{i}$ and $\mathbf{y}_{i}$, respectively. In this study, $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|);\quad m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$. At the end, the calculated control input $\mathbf{u}(k)$ needs to be denormalized before applying to the plant.
@@
8:  Denormalize the signal $\mathbf{u}_{N}^{*}(k)$
9:  Apply $u(k)=\mathbf{u}_{N_{[0:n_{u}-1]}}^{*}(k)$ to the system
@@ 4 Stability Analysis
This section establishes a rigorous stability analysis of the developed NTDPC approach. For both noise-free and noisy cases, we establish stability to the desired equilibrium. In the noise-free case, asymptotic stability ensures tracking of the reference. In the noisy case, input-to-state stability (ISS) guarantees robustness to bounded measurement noise. To facilitate the stability analysis, we define key concepts and notations as done in [24] for DeePC.
@@
Assume $Q,R$ are block-diagonal:
$\|\mathbf{y}_{N}(k)\|_{Q}^{2}$ $=\sum_{i=0}^{N-1}\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2},$
$\|\mathbf{u}_{N}(k)\|_{R}^{2}$ $=\sum_{i=0}^{N-1}\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2},$ (52)
where $Q_{i},R_{i}\succ 0$, and $r_{y}\in\mathbb{R}^{n_{y}}$, $r_{u}\in\mathbb{R}^{n_{u}}$ are the steady-state output and input, while $y^{*}(i|k),u^{*}(i|k)$ are the predicted output and input at step $i$ from time $k$. The stage cost is:
$l(y^{*}(i|k),u^{*}(i|k))$ $=\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2}.$
Thus, $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}$.
@@
The non-minimal state is:
$X_{\text{ini}}(k)$ $=\mathbf{z}_{\text{ini}}(k)=\text{col}(\mathbf{y}_{\text{ini}}(k),\mathbf{u}_{\text{ini}}(k))=\text{col}(y(k-T_{\text{ini}}),\dots,y(k-1),u(k-T_{\text{ini}}),\dots,u(k-1)),$
with $\mathbf{y}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}},\mathbf{u}_{\text{ini}}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}}$. The error state is:
$\tilde{X}_{\text{ini}}(k)$ $=\text{col}(\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y},\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u}),$
where $\mathbf{r}_{y}=\text{col}(r_{y},\dots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\dots,r_{u})$.
@@
A function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\varphi(0)=0$. It is $\mathcal{K}_{\infty}$ if $\lim_{s\to\infty}\varphi(s)=\infty$. A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for fixed $t\geq 0$, and $\beta(s,t)\to 0$ as $t\to\infty$ for fixed $s>0$. The identity is $\text{id}(s)=s$.
@@
The filtration $\mathcal{F}_{k}$ is the sigma-algebra generated by the measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available up to time $k$.

### v2

@@ 3.4 Dimension Reduction of the Future Trajectory Parameterization
We now unify the past and future parameterizations into a single low-order system. The past trajectory is given by $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Since $\mathbf{g}_{f}(k)$ lies in the null space of $\mathbf{Z}_{p}$, this simplifies to $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)$, which we previously derived in (17) as $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$.
For the future trajectory, we substitute $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ and the result from (30) into (28):
$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k).$ (31)
Defining the coupling matrix $S=\mathbf{Z}_{f}V_{1}$, we can express the full trajectory vector in terms of the latent variables $\eta_{1}(k)$ and $\eta_{2}(k)$:
$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$. (32)
To facilitate the predictor derivation, we partition the matrices $L_{1}$, $S$, and $L_{f_{1}}$ to separate input and output components:
$L_{1}=\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix},\quad S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$. (33)
where the dimensions of the partitioned matrices are defined consistent with the signal dimensions: $L_{u}\in\mathbb{R}^{n_{u}T_{ini}\times r}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{ini}\times r}$, $S_{u}\in\mathbb{R}^{n_{u}N\times r}$, $S_{y}\in\mathbb{R}^{n_{y}N\times r}$. This implicitly satisfies $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$.
@@ 3.5 Optimal Multi-Step Predictor
We apply the parameterization in (32) to construct an $N$-step ahead predictor. Our goal is to derive the minimum variance unbiased predictor of the future output $\mathbf{y}_{N}(k)$ given the estimated initial state and future control inputs. Expanding the lower block of (32):
$\begin{split}\mathbf{u}_{N}(k)&=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k),\\ \mathbf{y}_{N}(k)&=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k).\end{split}$ (34)
Substituting the black estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (34) yields the prediction model:
$\begin{split}\mathbf{u}_{N}(k)&=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k),\\ \mathbf{y}_{N}(k)&=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k).\end{split}$ (35)
Since $L_{f_{u}}$ is invertible, we can uniquely solve for the future latent vector $\eta_{2}(k)$ using the specified future inputs $\mathbf{u}_{N}(k)$:
@@
Substituting (36) into the output equation in (35), we obtain the explicit predictor $\hat{\mathbf{y}}_{N}(k)$:
@@
Replacing $\hat{\eta}_{1}(k)$ with the estimator $\mathcal{K}~\mathbf{z}_{\text{ini},m}(k)$ derived in the Theorem 3.1:
$\hat{\mathbf{y}}_{N}(k)=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathcal{K}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).$ (38)
We define the predictor system matrices as:
$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\mathcal{K},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (39)
where $P_{1}\in\mathbb{R}^{n_{y}N\times(n_{u}+n_{y})T_{ini}}$, $P_{2}\in\mathbb{R}^{n_{y}N\times n_{u}N}$. Then, it leads to the compact affine prediction model:
@@
To ensure the mathematical rigor of the NTDPC formulation, we introduce the following assumptions and lemma.
@@
[Future Excitation and Reduced-Order Parameterization] The future input sequence $u_{N}$ is sufficiently exciting such that
$\operatorname{rank}(Z_{f}V_{2})=n_{u}N,$
where $V_{2}$ spans the nullspace of $Z_{p}$. As a result, the reduced future latent variable $\eta_{2}\in\mathbb{R}^{n_{u}N}$ provides a complete parameterization of all admissible future trajectories consistent with the past data.
Remark 3.2.
The reduced future input mapping $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is assumed to be full rank in the noise-free case, which is guaranteed provided the offline training data satisfies the condition of persistency of excitation. In the presence of noise or numerical degeneracy, a regularized pseudoinverse of $L_{f_{u}}$ is employed to compute the future latent variable $\eta_{2}$. This guarantees the existence and uniqueness of the multi-step predictor in practice.
@@
The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex, closed, and non-empty.
@@
Under Assumptions 3.2 and invertible property of $L_{f_{u}}$, the predictor in (40) is an unbiased estimator of the true future output $\mathbf{y}_{N}^{\text{true}}(k)$ given the parameters $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$. That is, $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)\mid\eta_{1}(k),\mathbf{u}_{N}(k)]=\mathbf{y}_{N}^{\text{true}}(k)$.
@@
The true output of the system, given the latent state $\eta_{1}(k)$ and input $\mathbf{u}_{N}(k)$, is defined by the second line of (34): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$. Substituting the exact solution for $\eta_{2}(k)$, we get:
$\mathbf{y}_{N}^{\text{true}}(k)=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).$ (41)
Taking the expectation of the predictor (38), and noting that $\mathbf{u}_{N}(k)$ is deterministic and $\mathbb{E}[\hat{\eta}_{1}(k)\mid\eta_{1}(k)]=\eta_{1}(k)$ (from Theorem 3.1):
$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (42)
This confirms the predictor is unbiased. Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ rely only on the SVD of the offline data, they can be computed offline. ∎
We now formally state the NTDPC optimization problem.
Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC computes the optimal control sequence by solving the following convex optimization problem at each time step $k$:
@@
subject to $\mathbf{y}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (43)
$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y},$
where $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ represents the predicted future output used in the cost function, and $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable ensuring feasibility against noise and model mismatch. The weighting matrices $Q,R,\Lambda_{y}\succ 0$ are user-defined. The vector $\mathbf{z}_{\text{ini},m}(k)$ contains the most recent $T_{\text{ini}}$ input and noisy output measurements. Algorithm 1 summarizes the implementation steps.
@@
To prevent numerical bias during the SVD factorization of $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$, it is recommended to normalize the input and output data. This balances the singular values when inputs and outputs have significantly different magnitudes. We apply the scaling:
$\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\quad\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k),$ (44)
where $M_{u}=\text{diag}(m_{u_{1}},\dots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\dots,m_{y_{n_{y}}})$. In this study, the scaling factors are chosen as the mean absolute values of the signals: $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|)$ and $m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$. The computed optimal control input $\mathbf{u}_{N}^{*}(k)$ must be denormalized before being applied to the plant.
@@
8:Extract optimal first input $u^{*}(k)$ from $u_{N}^{*}(k)$
9:Denormalize $u^{*}(k)\leftarrow M_{u}u^{*}(k)$ and apply to system
@@ 4 Stability Analysis
This section establishes a rigorous stability analysis of the developed NTDPC approach. We examine the system’s behavior in both noise-free and noisy scenarios. In the noise-free case, we prove asymptotic stability, ensuring exact tracking of the reference. In the noisy case, we demonstrate Input-to-State Stability (ISS), guaranteeing robustness to bounded measurement noise. To facilitate this analysis, we define the following key concepts and notations, adapted from the dissipativity framework for data-driven control [22].
@@
Assume the weighting matrices $Q$ and $R$ are block-diagonal with blocks $Q_{i}\succ 0$ and $R_{i}\succ 0$. The stage cost $l(\cdot,\cdot)$ is defined as:
$l(\mathbf{y},\mathbf{u})=\|\mathbf{y}-r_{y}\|_{Q_{i}}^{2}+\|\mathbf{u}-r_{u}\|_{R_{i}}^{2},$ (45)
where $r_{y}\in\mathbb{R}^{n_{y}}$ and $r_{u}\in\mathbb{R}^{n_{u}}$ denote the steady-state output and input references, respectively. The total cost function over the prediction horizon $N$ is given by:
$J^{*}(k)=\sum_{i=0}^{N-1}l(\mathbf{y}^{*}(i|k),\mathbf{u}^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2},$ (46)
where $\mathbf{y}^{*}(i|k)$ and $\mathbf{u}^{*}(i|k)$ are the predicted output and input at step $i$ given information at time $k$.
@@
The non-minimal state vector $\mathbf{z}_{\text{ini}}(k)$ is defined as:
$\mathbf{z}_{\text{ini}}(k)=\operatorname{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)).$
The error state $\tilde{\mathbf{z}}_{\text{ini}}(k)$ represents the deviation from equilibrium:
$\tilde{\mathbf{z}}_{\text{ini}}(k)=\operatorname{col}(\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u},\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y}),$
where $\mathbf{r}_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}}$ and $\mathbf{r}_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ are stacked reference vectors.
@@
A continuous function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to class $\mathcal{K}$ if it is strictly increasing and $\varphi(0)=0$. It belongs to class $\mathcal{K}_{\infty}$ if it is also unbounded, i.e., $\lim_{s\to\infty}\varphi(s)=\infty$. A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to class $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for each fixed $t\geq 0$, and $\lim_{t\to\infty}\beta(s,t)=0$ for each fixed $s>0$. The identity function is denoted by $\operatorname{id}(s)=s$.
@@
The filtration $\mathcal{F}_{k}$ is the $\sigma$-algebra generated by the history of measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available to the controller up to time $k$.